Output AC. AFC. AVC. MC

1. 100. 40

2. 36

3. 36

4. 36

5. 12. 41


1
Expert's answer
2021-03-22T11:01:25-0400

Let's first find the fixed cost (FC):


"AFC_1=\dfrac{FC}{Q_1},""FC=AFC_1Q_1=\$40\cdot1=\$40."

Then, we can find "AFC" for the output from "Q=2" to "Q=5":


"AFC_2=\dfrac{FC}{Q_2}=\dfrac{\$40}{2}=\$20,""AFC_3=\dfrac{FC}{Q_3}=\dfrac{\$40}{3}=\$13.33,""AFC_4=\dfrac{FC}{Q_4}=\dfrac{\$40}{4}=\$10,""AFC_5=\dfrac{FC}{Q_5}=\dfrac{\$40}{5}=\$8."

Since, "AC=AFC+AVC", we can find "AVC" for each number of output:


"AVC_1=AC_1-AFC_1=\$100-\$40=\$60,""AVC_2=AC_2-AFC_2=\$36-\$20=\$16,""AVC_3=AC_3-AFC_3=\$36-\$13.33=\$22.67,""AVC_4=AC_4-AFC_4=\$36-\$10=\$26,""AVC_5=AC_5-AFC_5=\$12-\$8=\$4."

Let's find the TC. By the definition,


"AC=\dfrac{TC}{Q},""TC_1=AC_1\cdot Q_1=\$100\cdot1=\$100,""TC_2=AC_2\cdot Q_2=\$36\cdot2=\$72,""TC_3=AC_3\cdot Q_3=\$36\cdot3=\$108,""TC_4=AC_4\cdot Q_4=\$36\cdot4=\$144,""TC_5=AC_5\cdot Q_5=\$12\cdot5=\$60."

By the definition of MC we have:


"MC_1=\dfrac{TC_2-TC_1}{Q_2-Q_1}=\dfrac{\$72-\$100}{2-1}=-\$28,""MC_2=\dfrac{TC_3-TC_2}{Q_3-Q_2}=\dfrac{\$108-\$72}{3-2}=\$36,""MC_3=\dfrac{TC_4-TC_3}{Q_4-Q_3}=\dfrac{\$144-\$108}{4-3}=\$36,""MC_4=\dfrac{TC_5-TC_4}{Q_5-Q_4}=\dfrac{\$60-\$144}{5-4}=-\$84."

Finally, we get the following table: